OpenStudy (anonymous):

What is the maximum number of electrons that can be contained in the first, second, third, and fourth energy levels, respectively?

OpenStudy (xishem):

Each shell has a certain number of orbitals, or sub-shells, within it. These sub-shells are called s, p, d, and f. S can hold 2 electrons, p -- 6, d -- 10, and f -- 14. Shell one contains a single s orbital. It can hold 2 electrons. Shell two contains one s and one p orbital. It can hold 2 + 6 = 8 electrons. Shell three contains one s, one p, and one d orbital. It can hold 2 + 6 + 10 = 18 electrons. Shell four contains one s, one p, one d, and one f orbital. It can hold 2 + 6 + 10 + 14 = 32 electrons.

OpenStudy (anonymous):

\[2n ^{2}\] this equation is being followed for n=1. electrons will be 2 for n=2 , electrons will be 8 and so on..... n is the number of sub shell

OpenStudy (anonymous):

sorry n is number of shell